How much should you invest at 12% compounded monthly in order to purchase a $25,000 car in 5 years?
Answer by checkley77(12844)

x(1+.12/12)^5*12=25,000
x(1.01)^60=25,000
x*1.8167=25,000
x=25,000/1.8167
x=13,761 needs to be invested for 5 years.
Proof:
13,761(1+.12/12)^5*12=25,000
13,761(1.01)^60=25,000
13,761*1.8167=25,000
25,000~25,000